# Stablity of charged scalar hair on a charged black hole

### v1

@@ 4 Exact Solutions of the Field Equations
Figure 2 shows the radial profiles of the scalar field $\psi(r)$ for different values of mass $\mu$ and charge $q$ of the scalar field, with $V^{\prime}(r_{h})=1.0$ and $\psi(r_{h})=0.5$ at the horizon. All these profiles show a damped oscillatory behaviour with respect to the radial distance $r$. The profile with $\mu=0.1$ and $q=1.0$ shows a steep fall until $r\sim 2r_{h}$. We note that this steep fall in $\psi(r)$ is correlated with the non-monotonic behaviour of $f(r)$ and a steep fall of $\eta(r)$ as seen previously in Figure 1.
@@
The results shown in Figures 2 and 3 suggest possible existence of charged and massive scalar hair surrounding the charged black hole. Stability of such hairy solutions is discussed in Section 5.
@@
Figure 5: Radial profiles of the electrostatic potential $V(r)$ for different values of $V^{\prime}(r_{h})$ and scalar hair parameters $q$ and $\mu$, with the horizon condition $\psi(r_{h})=0.5$. For comparison, the profile of the Reisnner-Nordström potential is also shown.
Figure 5 shows the radial profiles of the electrostatic potential $V(r)$ with different values of the parameters $q$ and $\mu$ of the scalar field, with horizon conditions $V^{\prime}(r_{h})=1.0$, and $\psi(r_{h})=0.5$. All profiles of $V(r)$ show monotonic behaviours: the potential monotonically increases from zero at the horizon and appears to saturate to constant values at infinity. For the case $\mu$ fixed at $0.1$, as $q$ is increased, the profiles of $V(r)$ deviate further from the Reisnner-Nordstróm profile; $q$ changing from $0.7$ to $1.0$ shows a rapid deviation, reaching a value much higher the Reisnner-Nordström value at infinity. On the other hand, for the case $q$ fixed at $1.0$, the profiles undergo slight deviation as $\mu$ is increased from zero to $0.1$. It is obvious that the electrostatic potential $V(r)$ acquires additional contribution from the charged scalar hair so that it approaches a higher constant value at infinity.
@@ 5 Dynamical Stability Analyses
Consequently, we shall analyse the stability of the hairy solutions obtained in the previous sections. For definiteness, we shall carry out the stability analysis in two different ways. The first is based on Sturm-Liouville equation whereas the second is based on a Schrödinger-like equation for the dynamic perturbation.
@@
$\ddot{\tilde{\psi}}-\tilde{\gamma}^{2}\tilde{\psi}^{\prime\prime}$ $-\dot{\tilde{\psi}}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)-\left\{\frac{\tilde{f}^{\prime}}{\tilde{f}}-\tilde{\eta}^{\prime}+\frac{2}{r}\right\}\tilde{\gamma}^{2}\tilde{\psi}^{\prime}-q\tilde{\chi}\dot{\tilde{V}}+q\tilde{\chi}\tilde{V}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)$ (26)
$-2q\tilde{V}\dot{\tilde{\chi}}-q^{2}\tilde{V}^{2}\tilde{\psi}+\mu^{2}\tilde{f}e^{-2\tilde{\eta}}\tilde{\psi}=0$
@@
$\ddot{\tilde{\chi}}-\tilde{\gamma}^{2}\tilde{\chi}^{\prime\prime}$ $-\dot{\tilde{\chi}}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)-\left\{\frac{\tilde{f}^{\prime}}{\tilde{f}}-\tilde{\eta}^{\prime}+\frac{2}{r}\right\}\tilde{\gamma}^{2}\tilde{\chi}^{\prime}+q\tilde{\psi}\dot{\tilde{V}}-q\tilde{\psi}\tilde{V}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)$ (27)
$+2q\tilde{V}\dot{\tilde{\psi}}-q^{2}\tilde{V}^{2}\tilde{\chi}+\mu^{2}\tilde{f}e^{-2\tilde{\eta}}\tilde{\psi}=0.$
@@ 6 Discussion and Conclusion
We thus enquired whether an electrically charged scalar field surrounding a charged black hole can form a stable configuration around a charged black hole. This may be relevant in the astrophysical context when a black hole is surrounded by a charged plasma.
@@
Before proceeding any further, it was important to ascertain whether the necessary condition of all relevant energy conditions supplemented by the causality condition are obeyed by the Einstein-Maxwell-charged scalar field model. We therefore considered all energy conditions following from the mixed components of the energy-momentum tensor and the causality condition. Upon expressing the mixed components of the energy-momentum tensor in terms of the scalar field and other quantities, we found that all energy conditions coupled with the causality condition are satisfied. This suggests the possibility of existence of a charged hairy solution.
@@
Furthermore, the electric potential can take highly enhanced values compared to the normal Reisnner-Nordström case without a hair. This is due to additional contribution to the electric potential coming from the charged scalar hair in addition to that from the charged black hole.
@@
The fact that all energy conditions coupled with the causality condition are satisfied, in addition to satisfying the condition of stability of the static hairy solutions with respect to dynamical perturbations, suggest that a charged black hole can be surrounded by a stable and electrically charged scalar hair.

### v3

@@ 4.2 Numerical integration
The results shown in Figures 3 and 4 suggest possible existence of charged and massive scalar hair surrounding the charged black hole. Stability of such hairy solutions is discussed in Section 5.
@@
Figure 6 shows the radial profiles of the electrostatic potential $V(r)$ with different values of the parameters $q$ and $\mu$ of the scalar field, with horizon conditions $V^{\prime}(r_{h})=1.0$, and $\psi(r_{h})=0.5$. All profiles of $V(r)$ show monotonic behaviours: the potential monotonically increases from zero at the horizon and appears to saturate to constant values at infinity. For the case $\mu$ fixed at $0.1$, as $q$ is increased, the profiles of $V(r)$ deviate further from the Reissner-Nordström profile; $q$ changing from $0.7$ to $1.0$ shows a rapid deviation, reaching a value much higher the Reissner-Nordström value at infinity. On the other hand, for the case $q$ fixed at $1.0$, the profiles undergo slight deviation as $\mu$ is increased from zero to $0.1$. It is obvious that the electrostatic potential $V(r)$ acquires additional contribution from the charged scalar hair so that it approaches a higher constant value at infinity. Moreover, the magnitude of the electric field $V^{\prime}(r)$ vanishes as $r\rightarrow\infty$, as expected.
Figure 6: Radial profiles of the electrostatic potential $V(r)$ for different values of $V^{\prime}(r_{h})$ and scalar hair parameters $q$ and $\mu$, with the horizon condition $\psi(r_{h})=0.5$. For comparison, the profile of the Reissner-Nordström potential is also shown.
@@ 5 Dynamical stability analyses
Consequently, we shall analyse the stability of the hairy solutions obtained in the previous sections. For definiteness, we shall carry out the stability analysis in two different ways. The first is based on Sturm-Liouville equation [49, 50, 51, 52] whereas the second is based on a Schrödinger-like equation [9, 10] for the dynamic perturbation.
@@
$\ddot{\tilde{\psi}}-\tilde{f}^{2}e^{-2\tilde{\eta}}\tilde{\psi}^{\prime\prime}$ $-\dot{\tilde{\psi}}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)-\left\{\frac{\tilde{f}^{\prime}}{\tilde{f}}-\tilde{\eta}^{\prime}+\frac{2}{r}\right\}\tilde{f}^{2}e^{-2\tilde{\eta}}\tilde{\psi}^{\prime}-q\tilde{\chi}\dot{\tilde{V}}$ (26)
$+q\tilde{\chi}\tilde{V}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)-2q\tilde{V}\dot{\tilde{\chi}}-q^{2}\tilde{V}^{2}\tilde{\psi}+\mu^{2}\tilde{f}e^{-2\tilde{\eta}}\tilde{\psi}=0$
@@
$\ddot{\tilde{\chi}}-\tilde{f}^{2}e^{-2\tilde{\eta}}\tilde{\chi}^{\prime\prime}$ $-\dot{\tilde{\chi}}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)-\left\{\frac{\tilde{f}^{\prime}}{\tilde{f}}-\tilde{\eta}^{\prime}+\frac{2}{r}\right\}\tilde{f}^{2}e^{-2\tilde{\eta}}\tilde{\chi}^{\prime}+q\tilde{\psi}\dot{\tilde{V}}$ (27)
$-q\tilde{\psi}\tilde{V}\left(\frac{\dot{\tilde{f}}}{\tilde{f}}-\dot{\tilde{\eta}}\right)+2q\tilde{V}\dot{\tilde{\psi}}-q^{2}\tilde{V}^{2}\tilde{\chi}+\mu^{2}\tilde{f}e^{-2\tilde{\eta}}\tilde{\psi}=0.$
@@ 6 Discussion and conclusion
Subsequently, although charged hairy solutions were obtained in Mayo and Bekenstein’s model with polynomial scalar potential of sixth degree and the scalar charge much smaller than the black hole charge, a natural question arises as to whether a charged scalar hair would exist if the scalar field is taken to be a charged massive Klein-Gordon field (with quadratic potential) without any limitation on its charge.
@@
Before proceeding any further, it was important to ascertain whether the necessary condition (i) of all relevant energy conditions supplemented by the causality condition are obeyed by the Einstein-Maxwell-charged scalar field model. We therefore considered all energy conditions following from the mixed components of the energy-momentum tensor and the causality condition. Upon expressing the mixed components of the energy-momentum tensor in terms of the scalar field and other quantities, we found that all energy conditions coupled with the causality condition are satisfied. This suggests the possibility of existence of a charged hairy solution.
@@
Furthermore, the electric potential can take highly enhanced values compared to the normal Reissner-Nordström case without a hair. This is due to additional contribution to the electric potential coming from the charged scalar hair in addition to that from the charged black hole.
@@
We note in passing that in references [46, 47, 48] the potential of the scalar field was considered to be a polynomial of degree 6, with only even powers of $\phi$, having the form $V(\phi)=a\phi^{2}-b\phi^{4}+c\phi^{6}$. On the other hand, the potential in the present work is only quadratic, that is $V(\phi)=a\phi^{2}$. This difference in the forms of the potentials makes a crucial change in the solutions. In references [46, 47, 48], a non-oscillatory decaying solution was obtained in the polynomial model for the potential with specific choices for the parameters. On the other hand, we obtained a decaying oscillatory solution for the scalar hair with the quadratic potential. In addition, we confirmed the stability of our decaying oscillatory solutions.
The fact that all energy conditions coupled with the causality condition are satisfied, in addition to satisfying the condition of stability of the static hairy solutions with respect to dynamical perturbations, suggest that a Reissner-Nordström black hole can be surrounded by a stable and electrically charged scalar hair with quadratic scalar potential.
In this context, we note that our theory considers static hairy solutions and confirms their stability while it does not include the dynamics of how such hairy solutions form around the black hole. It is well-known that a Reissner-Nordström black hole may undergo the phenomenon of superradiance instability due to scattering of incident charged particles, giving rise to outgoing charges if the condition of superradiance is fulfilled [54]. However, for an asymptotically flat Reissner-Nordström spacetime, the necessary conditions required to trigger a possible superradiance instability cannot be satisfied [55, 56]. Thus the origin of charged scalar hair in our model does not lie in the phenomenon of superradiance instability and it remains a matter of exploration in the future.